# Cointegration with Occasionally Binding Constraints

### v3

@@ 1 Introduction
As in a linear VAR, unit roots are unavoidable if one wishes to apply the CKSVAR to series with stochastic trends. The usual criticism of simply estimating a model in differences – that this obliterates the identifying long-run information carried by the cointegrating relations – is here magnified by the threshold nonlinearity in the model, which dictates whether the affected variable (e.g. interest rates) should enter in levels or differences, and in turn prescribes appropriate forms for the other variables. [^1] That these other variables (which enter the model linearly) cannot generally be replaced by their first differences can be seen from the vector error correction form of the CKSVAR given in (LABEL:vecm-general) below, from which it is evident that making this replacement (of $x_{t}$ by $\Delta x_{t}$) would amount to imposing the restriction that $\Pi^{x}=0$. To put it another way, that nonlinearity prevents series from being simply ‘differenced to stationarity’. Moreover, we show that in our setting, unit roots are not a mere technical nuisance: rather, their presence may impart significant identifying power to the low frequency behaviour of the series. For instance, the possibility that cointegrating relations between series may change as one series crosses a threshold, something accommodated by the CKSVAR, may be utilised to test hypotheses about the relative effectiveness of unconventional monetary policy (see Examples 2.1 and 2.1 below; this is a problem that has been studied econometrically by e.g. Gambacorta et al., 2014; Wu and Xia, 2016; Debortoli et al., 2020; and Ikeda et al., 2024).
@@
In analysing the CKSVAR with unit roots, we make a major contribution to the literature on nonlinear cointegration. Here a long-standing open problem relates to how a (nonlinear) vector autoregression, which provides a unified description of the short- and long-run dynamics of a collection of time series, can generate nonlinear cointegration between those series – where ‘nonlinear cointegration’ is understood in the profound sense of those series having common nonlinear stochastic trends with possibly nonlinear cointegrating relations between those trends. As discussed in the recent review by Tjøstheim (2020), despite the voluminous literature on the subject of ‘nonlinear cointegration’, this problem has yet to be addressed at any reasonable level of generality. [^2] While Cai et al. (2017) make an important effort in this direction, their results are limited to a first-order bivariate VAR with two regimes, in which one of those regimes is delimited by a compact set, and so makes a negligible contribution to the long-run behaviour of the series generated by the model. Their results are thus markedly different from those obtained below. Within the framework of the CKSVAR, we provide a resolution of this problem, showing that the model naturally gives rise to nonlinear cointegration, generating nonlinear common trend processes – censored, regulated, and kinked Brownian motions (see 3.3 below) – that have not previously appeared in multivariate settings.
@@
In developing a VAR model that exhibits both nonlinearity in its short-run dynamics, as in (LABEL:nlvecm), and in the implied (long-run) common trends, as in (LABEL:nlcointreg), this paper is the first to bridge the remarkably wide gulf that has existed between these two strands of the literature. The CKSVAR turns out to provide just enough flexibility to accommodate meaningful departures from linear cointegration, while retaining a sufficiently nice structure to be tractable. We show that depending on the rank conditions imposed on (submatrices of) the autoregressive polynomial evaluated at unity, the CKSVAR is capable of generating three distinct kinds of nonlinear cointegration, which we term: (i) regulated cointegration; (ii) kinked cointegration; and (iii) linear cointegration in a nonlinear VECM.
At a technical level, our contribution consists of identifying the alternative configurations of the model that give rise to cases (i)–(iii), which are essentially exhaustive of the possibilities here, and deriving analogues of the Granger–Johansen representation theorem in these three cases. [^3] Our analysis is exhaustive with respect to the possibilities for generating series that are integrated of order one (in a suitably extended sense of the term; see 3.1) within the CKSVAR; higher orders of integration are not considered here. In analysing case (i), our work relates to that of Cavaliere (2005), Liu et al. (2011), Gao et al. (2013), and most closely to Bykhovskaya and Duffy (2024), all of whom obtain convergence to regulated Brownian motions in univariate models. Case (ii), even in the univariate setting, does not appear to have been considered by any previous literature. While cases (i) and (ii) describe phenomena that are entirely new to the literature, case (iii) holds under a configuration of the model that falls within the very broad class of nonlinear VECMs considered by Saikkonen (2008): but even here, our results regarding the ergodicity of the short-memory components of the model extend his, insofar as we are able to exploit certain properties of the CKSVAR that are not shared by all the models encompassed by his general framework.
@@
The remainder of the paper is organised as follows. 2 introduces the CKSVAR model and the stylised structural macroeconomic models that we use as running examples. 3 develops the heuristics of the model with unit roots, outlining the tripartite classification noted above. The main results of this paper, which extend the Granger–Johansen representation theorem to the CKSVAR model, are given in 4. All proofs appear in the appendices.
@@
[…]
@@ 2 Model: the censored and kinked SVAR
We consider a VAR($k$) model in $p$ variables, in which one series, $y_{t}$, enters with coefficients that differ according to whether it is above or below a time-invariant threshold $b$, while the other $p-1$ series, collected in $x_{t}$, enter linearly. Defining
@@
for $\phi_{i}^{\pm}\in\mathbb{R}^{p\times 1}$ and $\Phi_{i}^{x}\in\mathbb{R}^{p\times(p-1)}$, and $L$ denotes the lag operator. As in a linear SVAR, $\{u_{t}\}$ may be an i.i.d. sequence of mutually orthogonal structural shocks, but our results below also permit them to be cross-correlated or weakly dependent. Through an appropriate redefinition of $y_{t}$ and $c$, we may take $b=0$ without loss of generality, and will do so throughout the sequel. [^4] Defining $y_{b,t}\coloneqq y_{t}-b$, $y_{b,t}^{+}\coloneqq\max\{y_{b,t},0\}$, $y_{b,t}^{-}\coloneqq\min\{y_{b,t},0\}$ and $c_{b}\coloneqq c-[\phi^{+}(1)+\phi^{-}(1)]b$, we can rewrite (LABEL:var-pm) as $\phi^{+}(L)y_{b,t}^{+}+\phi^{-}(L)y_{b,t}^{-}+\Phi^{x}(L)x_{t}=c_{b}+u_{t}.$ In this case, $y_{t}^{+}$ and $y_{t}^{-}$ respectively equal the positive and negative parts of $y_{t}$, and $y_{t}=y_{t}^{+}+y_{t}^{-}$. (Throughout the following, the notation ‘$a^{\pm}$’ connotes $a^{+}$ and $a^{-}$ as objects associated respectively with $y_{t}^{+}$ and $y_{t}^{-}$, or their lags. If we want to instead denote the positive and negative parts of some $a\in\mathbb{R}$, we shall do so by writing $[a]_{+}\coloneqq\max\{a,0\}$ or $[a]_{-}\coloneqq\min\{a,0\}$.)
@@
Models of the form of (LABEL:var-pm) have previously been employed in the literature to account for the dynamic effects of censoring, occasionally binding constraints, and endogenous regime switching. Mavroeidis (2021) proposed exactly this model, which he termed the censored and kinked structural VAR (CKSVAR) model, to describe the operation of monetary policy during periods when a zero lower bound may bind on the policy rate: in our notation, $y_{t}$ corresponds to his ‘shadow rate’, expressing the central bank’s desired policy stance, and $y_{t}^{+}$ to the actual policy rate. Aruoba et al. (2022) considered a model in which one variable is subject to an occasionally binding constraint, which although in its initial formulation is somewhat more general, reduces to an instance of the CKSVAR once the conditions necessary for the model to have a unique solution (for all values of $u_{t}$) have been imposed (see their Proposition 1(i)). This version of their model – i.e. that in which the ‘private sector regression functions’ are piecewise linear and continuous – is thus accommodated by (LABEL:var-pm). [^5] See also Aruoba et al. (2021), for a DSGE model with an occasionally binding constraint, in which agents’ decision rules are approximated by functions with these properties.
@@
To put some economic flesh on the bones of the representation theory developed in Sections 3 and 4 below, we here introduce the running example of a stylised structural model of monetary policy in the presence of a zero lower bound (ZLB). This model provides a simple, economically interpretable framework in which we may illustrate the various forms of novel long-run behaviour permitted by the CKSVAR, by considering alternate parametrisations of the model. Moreover, as discussed in 3 below, the model elucidates how that long-run behaviour may provide identifying information on the relative effectiveness of unconventional monetary policy (as compared with conventional rate-setting policy), i.e. on whether the zero lower bound really constrains the ability of a central bank to target inflation.
@@
To ‘close’ the model, we consider two alternative specifications for the underlying processes followed by $\{r_{t}^{\ast}\}$ and $\{\overline{\pi}_{t}\}$. In the first of these (henceforth, Example 2.1a) the inflation target is assumed to be constant and is normalised to zero (i.e. $\overline{\pi}_{t}=\overline{\pi}=0$), while the natural real rate of interest follows an AR(1) process,
$r_{t}^{\ast}=\psi r_{t-1}^{\ast}+\eta_{t}$ (2.6)
where $\psi\in(-1,1]$, and $\eta_{t}$ is an i.i.d. mean zero innovation, possibly correlated with $\varepsilon_{t}$. (When $\psi=1$, $\{r_{t}^{\ast}\}$ is integrated, as in the model of Laubach and Williams, 2003.) Substituting (LABEL:taylor) into (LABEL:is-pc) and (LABEL:natural-rate), we render the system as a CKSVAR for $(i_{t},\pi_{t})$ as
$\begin{bmatrix}1&1&-\gamma\\ 0&\theta(1-\mu)&1-\theta\gamma\end{bmatrix}\begin{bmatrix}i_{t}^{+}\\ i_{t}^{-}\\ \pi_{t}\end{bmatrix}=\begin{bmatrix}\psi&\psi&-\psi\gamma\\ 0&0&0\end{bmatrix}\begin{bmatrix}i_{t-1}^{+}\\ i_{t-1}^{-}\\ \pi_{t-1}\end{bmatrix}+\begin{bmatrix}\eta_{t}\\ \varepsilon_{t}\end{bmatrix}.$ (2.7)
This model will provide an illustration of the second kind of nonlinear cointegration developed in 3 below.
@@
where $\delta\in(-1,0]$, and $\eta_{t}$ is an i.i.d. innovation as above. When $\delta=0$, this corresponds to a model in which the inflation target follows a pure random walk, possibly reflecting the time-varying preferences of the central bank (cf. Cogley and Sbordone, 2008); when $\delta<0$, the model allows past deviations of inflation from target to feed back into the target, such that e.g. below-target inflation induces an upward revision of the inflation target. Motivation for this aspect of the model comes from the manner in which the ZLB may constrain policy to be excessively deflationary for a sustained period, something that has prompted the literature to consider the costs and benefits of adopting a higher inflation target (e.g. Blanchard et al., 2010, pp. 207f.; Coibion et al., 2012). Supposing additionally that $\gamma>1$, we may put (LABEL:is-pc), (LABEL:taylor) and (LABEL:infltarget) in the form of a CKSVAR as
@@
where $\varphi_{\mu}\coloneqq(1-\mu\theta\gamma)-\theta(1-\mu)$ and so $\varphi_{1}=1-\theta\gamma$. Depending on the assumptions made on the model parameters (in particular $\delta$), this model is capable of generating either of the first two types of nonlinear cointegration discussed in 3.
While both Mavroeidis (2021) and Aruoba et al. (2022) motivate and interpret (LABEL:var-pm) as a structural model, empirically motivated reduced-form models of this kind have also appeared in the literature, particularly in the univariate ($p=1$) case of (LABEL:var-pm), which encompasses the dynamic Tobit model (Maddala 1983, p. 186; for applications, see e.g. Demiralp and Jordà, 2002; De Jong and Herrera, 2011; Brezigar-Masten et al., 2021; and Bykhovskaya, 2023).
@@
Consider (LABEL:var-pm) with $p=1$ and $\phi_{0}^{+}=\phi_{0}^{-}=1$, so that
@@
We follow Mavroeidis (2021) and Aruoba et al. (2022) in maintaining the following, which are necessary and sufficient to ensure that (LABEL:var-pm) has a unique solution for $(y_{t},x_{t})$, for all possible values of $u_{t}$. Define
@@
    $\{(y_{t},x_{t})\}$ are generated according to (LABEL:y-threshold)–(LABEL:var-pm) with $b=0$, with (possibly random) initial values $(y_{i},x_{i})$, for $i\in\{-k+1,\ldots,0\}$;
@@
While it is not always the case that the reduced form of (LABEL:var-pm) corresponds directly to a canonical CKSVAR, by defining the canonical variables
@@
Suppose DGP holds. Then there exist $(\tilde{y}_{t},\tilde{x}_{t})$ such that (LABEL:canon-vars)–(LABEL:canon-polys) hold, $\tilde{y}_{t}^{+}=\max\{\tilde{y}_{t},0\}$, $\tilde{y}_{t}^{-}=\min\{\tilde{y}_{t},0\}$ and
@@
To distinguish between a general CKSVAR in which possibly $\Phi_{0}\neq I_{p}^{\ast}$, and its associated canonical form as given by 2.1, we shall refer to the former as the ‘structural form’ of the CKSVAR. Since the time series properties of a general CKSVAR are largely inherited from its derived canonical form, we shall often work with this more convenient representation of the system, and indicate this as follows.
@@
$\{(y_{t},x_{t})\}$ are generated by a canonical CKSVAR, i.e. DGP holds with $\Phi_{0}=[\phi_{0}^{+},\phi_{0}^{-},\Phi^{x}]=I_{p}^{\ast}$, so that (LABEL:var-two-sided) may be equivalently written as

### v4

@@ 1 Introduction
As in a linear VAR, unit roots are unavoidable if one wishes to apply the CKSVAR to series with stochastic trends. The usual criticism of simply estimating a model in differences – that this obliterates the identifying long-run information carried by the cointegrating relations – is here magnified by the threshold nonlinearity in the model, which dictates whether the affected variable (e.g. interest rates) should enter in levels or differences, and in turn prescribes appropriate forms for the other variables. [^1] That these other variables (which enter the model linearly) cannot generally be replaced by their first differences can be seen from the vector error correction form of the CKSVAR given in (3.1) below, from which it is evident that making this replacement (of $x_{t}$ by $\Delta x_{t}$) would amount to imposing the restriction that $\Pi^{x}=0$. To put it another way, that nonlinearity prevents series from being simply ‘differenced to stationarity’. Moreover, we show that in our setting, unit roots are not a mere technical nuisance: rather, their presence may impart significant identifying power to the low frequency behaviour of the series. For instance, the possibility that cointegrating relations between series may change as one series crosses a threshold, something accommodated by the CKSVAR, may be utilised to test hypotheses about the relative effectiveness of unconventional monetary policy (see Examples 2.1 and 2.1 below; this is a problem that has been studied econometrically by e.g. Gambacorta et al., 2014; Wu and Xia, 2016; Debortoli et al., 2020; and Ikeda et al., 2024).
@@
In analysing the CKSVAR with unit roots, we make a major contribution to the literature on nonlinear cointegration. Here a long-standing open problem relates to how a (nonlinear) vector autoregression, which provides a unified description of the short- and long-run dynamics of a collection of time series, can generate nonlinear cointegration between those series – where ‘nonlinear cointegration’ is understood in the profound sense of those series having common nonlinear stochastic trends with possibly nonlinear cointegrating relations between those trends. As discussed in the recent review by Tjøstheim (2020), despite the voluminous literature on the subject of ‘nonlinear cointegration’, this problem has yet to be addressed at any reasonable level of generality. [^2] While Cai et al. (2017) make an important effort in this direction, their results are limited to a first-order bivariate VAR with two regimes, in which one of those regimes is delimited by a compact set, and so makes a negligible contribution to the long-run behaviour of the series generated by the model. Their results are thus markedly different from those obtained below. Within the framework of the CKSVAR, we provide a resolution of this problem, showing that the model naturally gives rise to nonlinear cointegration, generating nonlinear common trend processes – censored, regulated, and kinked Brownian motions (see Definition 3.3 below) – that have not previously appeared in multivariate settings.
@@
In developing a VAR model that exhibits both nonlinearity in its short-run dynamics, as in (1.1), and in the implied (long-run) common trends, as in (1.2), this paper is the first to bridge the remarkably wide gulf that has existed between these two strands of the literature. The CKSVAR turns out to provide just enough flexibility to accommodate meaningful departures from linear cointegration, while retaining a sufficiently nice structure to be tractable. We show that depending on the rank conditions imposed on (submatrices of) the autoregressive polynomial evaluated at unity, the CKSVAR is capable of generating three distinct kinds of nonlinear cointegration, which we term: (i) regulated cointegration; (ii) kinked cointegration; and (iii) linear cointegration in a nonlinear VECM.
At a technical level, our contribution consists of identifying the alternative configurations of the model that give rise to cases (i)–(iii), which are essentially exhaustive of the possibilities here, and deriving analogues of the Granger–Johansen representation theorem in these three cases. [^3] Our analysis is exhaustive with respect to the possibilities for generating series that are integrated of order one (in a suitably extended sense of the term; see Definition 3.1) within the CKSVAR; higher orders of integration are not considered here. In analysing case (i), our work relates to that of Cavaliere (2005), Liu et al. (2011), Gao et al. (2013), and most closely to Bykhovskaya and Duffy (2024), all of whom obtain convergence to regulated Brownian motions in univariate models. Case (ii), even in the univariate setting, does not appear to have been considered by any previous literature. While cases (i) and (ii) describe phenomena that are entirely new to the literature, case (iii) holds under a configuration of the model that falls within the very broad class of nonlinear VECMs considered by Saikkonen (2008): but even here, our results regarding the ergodicity of the short-memory components of the model extend his, insofar as we are able to exploit certain properties of the CKSVAR that are not shared by all the models encompassed by his general framework.
@@
The remainder of the paper is organised as follows. Section 2 introduces the CKSVAR model and the stylised structural macroeconomic models that we use as running examples. Section 3 develops the heuristics of the model with unit roots, outlining the tripartite classification noted above. The main results of this paper, which extend the Granger–Johansen representation theorem to the CKSVAR model, are given in Section 4. All proofs appear in the appendices.
@@
$e_{m,i}$ denotes the $i$th column of an $m\times m$ identity matrix; when $m$ is clear from the context, we write this simply as $e_{i}$. In a statement such as $f(a^{\pm},b^{\pm})=0$, the notation $\pm$’ signifies that both $f(a^{+},b^{+})=0$ and $f(a^{-},b^{-})=0$ hold; similarly, $a^{\pm}\in A$’ denotes that both $a^{+}$ and $a^{-}$ are elements of $A$. All limits are taken as $n\rightarrow\infty$ unless otherwise stated. $\overset{p}{\rightarrow}$ and $\rightsquigarrow$ respectively denote convergence in probability and in distribution (weak convergence). We write $X_{n}(\lambda)\rightsquigarrow X(\lambda)$ on $D_{\mathbb{R}^{m}}[0,1]$’ to denote that $\{X_{n}\}$ converges weakly to $X$, where these are considered as random elements of $D_{\mathbb{R}^{m}}[0,1]$, the space of cadlag functions $0,1]\rightarrow\mathbb{R}^{m}$, equipped with the uniform topology; we denote this as $D[0,1]$ whenever the value of $m$ is clear from the context. $\lVert\cdot\rVert$ denotes the Euclidean norm on $\mathbb{R}^{m}$, and the matrix norm that it induces. For $X$ a random vector and $p\geq 1$, $\lVert X\rVert_{p}\coloneqq(\mathbb{E}\lVert X\rVert^{p})^{1/p}$.
@@ 2 Model: the censored and kinked SVAR
We consider a VAR($k$) model in $p$ variables, in which one series, $y_{t}$, enters with coefficients that differ according to whether it is above or below a time-invariant threshold $b$, while the other p1p-1 series, collected in $x_{t}$, enter linearly. Defining
@@
for $\phi_{i}^{\pm}\in\mathbb{R}^{p\times 1}$ and $\Phi_{i}^{x}\in\mathbb{R}^{p\times(p-1)}$, and $L$ denotes the lag operator. As in a linear SVAR, $\{u_{t}\}$ may be an i.i.d. sequence of mutually orthogonal structural shocks, but our results below also permit them to be cross-correlated or weakly dependent. Through an appropriate redefinition of $y_{t}$ and $c$, we may take $b=0$ without loss of generality, and will do so throughout the sequel. [^4] Defining $y_{b,t}\coloneqq y_{t}-b$, $y_{b,t}^{+}\coloneqq\max\{y_{b,t},0\}$, $y_{b,t}^{-}\coloneqq\min\{y_{b,t},0\}$ and $c_{b}\coloneqq c-[\phi^{+}(1)+\phi^{-}(1)]b$, we can rewrite (2.3) as $\phi^{+}(L)y_{b,t}^{+}+\phi^{-}(L)y_{b,t}^{-}+\Phi^{x}(L)x_{t}=c_{b}+u_{t}.$ In this case, $y_{t}^{+}$ and $y_{t}^{-}$ respectively equal the positive and negative parts of $y_{t}$, and $y_{t}=y_{t}^{+}+y_{t}^{-}$. (Throughout the following, the notation $a^{\pm}$’ connotes $a^{+}$ and $a^{-}$ as objects associated respectively with $y_{t}^{+}$ and $y_{t}^{-}$, or their lags. If we want to instead denote the positive and negative parts of some $a\in\mathbb{R}$, we shall do so by writing $a]_{+}\coloneqq\max\{a,0\}$ or $a]_{-}\coloneqq\min\{a,0\}$.)
@@
Models of the form of (2.3) have previously been employed in the literature to account for the dynamic effects of censoring, occasionally binding constraints, and endogenous regime switching. Mavroeidis (2021) proposed exactly this model, which he termed the censored and kinked structural VAR (CKSVAR) model, to describe the operation of monetary policy during periods when a zero lower bound may bind on the policy rate: in our notation, $y_{t}$ corresponds to his ‘shadow rate’, expressing the central bank’s desired policy stance, and $y_{t}^{+}$ to the actual policy rate. Aruoba et al. (2022) considered a model in which one variable is subject to an occasionally binding constraint, which although in its initial formulation is somewhat more general, reduces to an instance of the CKSVAR once the conditions necessary for the model to have a unique solution (for all values of $u_{t}$) have been imposed (see their Proposition 1(i)). This version of their model – i.e. that in which the ‘private sector regression functions’ are piecewise linear and continuous – is thus accommodated by (2.3). [^5] See also Aruoba et al. (2021), for a DSGE model with an occasionally binding constraint, in which agents’ decision rules are approximated by functions with these properties.
@@
To put some economic flesh on the bones of the representation theory developed in Sections 3 and 4 below, we here introduce the running example of a stylised structural model of monetary policy in the presence of a zero lower bound (ZLB). This model provides a simple, economically interpretable framework in which we may illustrate the various forms of novel long-run behaviour permitted by the CKSVAR, by considering alternate parametrisations of the model. Moreover, as discussed in Section 3 below, the model elucidates how that long-run behaviour may provide identifying information on the relative effectiveness of unconventional monetary policy (as compared with conventional rate-setting policy), i.e. on whether the zero lower bound really constrains the ability of a central bank to target inflation.
@@
To ‘close’ the model, we consider two alternative specifications for the underlying processes followed by $\{r_{t}^{\ast}\}$ and $\{\overline{\pi}_{t}\}$. In the first of these (henceforth, Example 2.1a) the inflation target is assumed to be constant and is normalised to zero (i.e. $\overline{\pi}_{t}=\overline{\pi}=0$), while the natural real rate of interest follows a random walk AR(1) process (as in the model of Laubach and Williams, 2003),
$r_{t}^{\ast}=r_{t-1}^{\ast}+\eta_{t}$ (2.6)
where $\eta_{t}$ is an i.i.d. mean zero innovation, possibly correlated with $\varepsilon_{t}$. Substituting (2.5) into (2.4) and (2.6), we render the system as a CKSVAR for $(i_{t},\pi_{t})$ as
$\begin{bmatrix}1&1&-\gamma\\ 0&\theta(1-\mu)&1-\theta\gamma\end{bmatrix}\begin{bmatrix}i_{t}^{+}\\ i_{t}^{-}\\ \pi_{t}\end{bmatrix}=\begin{bmatrix}1&1&-\gamma\\ 0&0&0\end{bmatrix}\begin{bmatrix}i_{t-1}^{+}\\ i_{t-1}^{-}\\ \pi_{t-1}\end{bmatrix}+\begin{bmatrix}\eta_{t}\\ \varepsilon_{t}\end{bmatrix}$. (2.7)
This model will provide an illustration of the second kind of nonlinear cointegration developed in Section 3 below.
@@
where $\delta\in(-1,0]$, and $\eta_{t}$ is an i.i.d. innovation as above. When $\delta=0$, this corresponds to a model in which the inflation target follows a pure random walk, possibly reflecting the time-varying preferences of the central bank (cf. Cogley and Sbordone, 2008); when $\delta<0$, the model allows past deviations of inflation from target to feed back into the target, such that e.g. below-target inflation induces an upward revision of the inflation target. Motivation for this aspect of the model comes from the manner in which the ZLB may constrain policy to be excessively deflationary for a sustained period, something that has prompted the literature to consider the costs and benefits of adopting a higher inflation target (e.g. Blanchard et al., 2010, pp. 207f.; Coibion et al., 2012). Supposing additionally that $\gamma>1$, we may put (2.4), (2.5) and (2.8) in the form of a CKSVAR as
@@
where $\varphi_{\mu}\coloneqq(1-\mu\theta\gamma)-\theta(1-\mu)$ and so $\varphi_{1}=1-\theta\gamma$. Depending on the assumptions made on the model parameters (in particular $\delta$), this model is capable of generating either of the first two types of nonlinear cointegration discussed in Section 3.
While both Mavroeidis (2021) and Aruoba et al. (2022) motivate and interpret (2.3) as a structural model, empirically motivated reduced-form models of this kind have also appeared in the literature, particularly in the univariate ($p=1$) case of (2.3), which encompasses the dynamic Tobit model (Maddala 1983, p. 186; for applications, see e.g. Demiralp and Jordà, 2002; De Jong and Herrera, 2011; Brezigar-Masten et al., 2021; and Bykhovskaya, 2023).
@@
Consider (2.3) with $p=1$ and $\phi_{0}^{+}=\phi_{0}^{-}=1$, so that
@@
We follow Mavroeidis (2021) and Aruoba et al. (2022) in maintaining the following, which are necessary and sufficient to ensure that (2.3) has a unique solution for $(y_{t},x_{t})$, for all possible values of $u_{t}$. Define
@@
    $\{(y_{t},x_{t})\}$ are generated according to (2.1)–(2.3) with $b=0$, with (possibly random) initial values $(y_{i},x_{i})$, for $i\in\{-k+1,\ldots,0\}$;
@@
While it is not always the case that the reduced form of (2.3) corresponds directly to a canonical CKSVAR, by defining the canonical variables
@@
Suppose DGP holds. Then there exist $(\tilde{y}_{t},\tilde{x}_{t})$ such that (2.13)–(2.14) hold, $\tilde{y}_{t}^{+}=\max\{\tilde{y}_{t},0\}$, $\tilde{y}_{t}^{-}=\min\{\tilde{y}_{t},0\}$ and
@@
To distinguish between a general CKSVAR in which possibly $\Phi_{0}\neq I_{p}^{\ast}$, and its associated canonical form as given by Proposition 2.1, we shall refer to the former as the ‘structural form’ of the CKSVAR. Since the time series properties of a general CKSVAR are largely inherited from its derived canonical form, we shall often work with this more convenient representation of the system, and indicate this as follows.
@@
$\{(y_{t},x_{t})\}$ are generated by a canonical CKSVAR, i.e. DGP holds with $\Phi_{0}=[\phi_{0}^{+},\phi_{0}^{-},\Phi^{x}]=I_{p}^{\ast}$, so that (2.2) may be equivalently written as